Technical Staff Blog

Author archives: John Bazik

RSS feed of John Bazik

Last update on .

The Technical Staff has several new members.  Paul Vars joined tstaff back in June as Senior Hardware Technician.  Shaun Wallace became Senior Systems Programmer last month, and David Serpa takes on the role of Lead Systems Administrator today.  With David's arrival, the technical staff has a full roster for the first time since 2016 ...

Last update on .

On the morning of Wednesday, January 17th, CIS will change the VLAN numbers of seven of the department's virtual network segments, affecting up to 422 systems, mostly desktop machines.  A simple bookkeeping change, users who are up very early that morning should notice only a brief network outtage.   The change is necessary for a ...

Last update on .

Due to campus-wide service outages, the main sections of the CS Dept website, including the home page, have been inaccessible for an extended period of time today, and remain down.  We are monitoring CIS' progress in addressing their issues, and we will ensure that the CS website returns to service when they are resolved.

Last update on .

For several days in the past week email messages sent to some CS users were forwarded to an incorrect address, causing them to "bounce."  Those messages were not delivered, and the senders were informed that the intended recipient's account does not exist.  This was due to an error on our part that has been ...

Last update on .

This Thursday, August 24th, your CS login password will change.  From that day forward, to log into a CS Linux system, you will use your Brown University login credentials.  For those of us whose "short id" (e.g. jbazik) is different from our CS logname (e.g. jsb), we will thenceforth use our short id ...

Last update on .

Today's service downtime for mailing lists, time machine and ownCloud due to filesystem maintenance will be extended until noontime.  This post will be updated when service returns.

UPDATE: CIS restored file services a bit before 2pm today.  Mailing list services were restored at around 2:30pm.

UPDATE​​​​​​​: Time Mchine has been restored.

Last update on .

Two GPFS NFS servers failed simultaneously at around 1pm today, causing file services to be unavailable for about 40 minutes.  The servers, crows and runts, serve different cluster groups; crows serves the grid and runts serves the internal department network.  Runts' failure was a kernel lock-up which prevented the normal automatic failover behavior.  The technical ...